APPLICATION:





an experimental study was conducted by a researcher to





determine if a new time slot has an effect on the performance





of pupils in mathematics. Fifteen randomly selected learners





participated in the study. Toward the end of the investigation, a





standardized assessment was conducted. The sample mean x̄= 75 and s=5. In standardization of the test, the mean was





65 and standard deviation was 8. Based on the evidence at





hand, is the new time slot effective? Use α = 0.05.




1
Expert's answer
2022-05-05T16:53:09-0400

Difference of two independent normal variables

Let "X" have a normal distribution with mean "\mu_X" and variance "\sigma_X^2."

Let "Y" have a normal distribution with mean "\mu_Y" and variance "\sigma_Y^2."

If "X" and "Y"are independent, then "X-Y"will follow a normal distribution with mean "\mu_X-\mu_Y" and variance "\sigma_X^2+\sigma_Y^2."

"\mu_{X-Y}=75-65=10""s_{X-Y}=\sqrt{(5)^2+(8)^2}=\sqrt{89}"

The following null and alternative hypotheses need to be tested:

"H_0:\mu=0"

"H_1:\mu\not=0"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.05,"

"df=n-1=15-1=14" degrees of freedom, and the critical value for a two-tailed test is "t_c=2.144787."

The rejection region for this two-tailed test is "R=\{t:|t|>2.144787\}"

The t-statistic is computed as follows:


"t=\dfrac{\mu_{X-Y}-\mu}{s_{X-Y}/\sqrt{n}}=\dfrac{10-0}{\sqrt{89}/\sqrt{15}}\approx4.105354"

Since it is observed that "|t|=4.105354>2.144787=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed "\alpha=0.05," "df=14, t=4.105354" is "p=0.001071," and since "p=0.001071<0.05=\alpha," it is then concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 0, at the "\alpha=0.05" significance level.

Therefore, there is enough evidence to claim that the new time slot can be effective, at the "\alpha=0.05" significance level.